OpenStudy (anonymous):

can some one help me with 7 rational expressions?

OpenStudy (anonymous):

\[\frac{ x+2 }{ (x-4)^{2} }- \frac{ x }{x-4 }\] this is my first one

OpenStudy (anonymous):

hey @Hero can you help me out?

hero (hero):

Hint: Multiply the second fraction by (x-4)/(x-4)

OpenStudy (anonymous):

\[\frac{ x^2-4 }{ x^2-16 }?\]

OpenStudy (anonymous):

@Hero is this right?

hero (hero):

No, it's not

OpenStudy (anonymous):

how would i go about doing this?

hero (hero):

I'm only going to show you this once.

hero (hero):

\[\space\space\space\space\frac{x + 2}{(x - 4)^2} - \frac{x}{x-4}\]\[=\frac{x + 2}{(x - 4)^2} - \frac{x}{x-4} \times\frac{x-4}{x-4}\]\[=\frac{x + 2}{(x - 4)^2} - \frac{x(x-4)}{x-4(x-4)}\]\[=\frac{x + 2}{(x - 4)^2} - \frac{x^2-4x}{(x-4)^2}\]\[=\frac{(x + 2) - (x^2 - 4x)}{(x-4)^2}\]\[=\frac{x + 2 - x^2 + 4x}{(x-4)^2}\]\[=\frac{-x^2 + 4x + x + 2}{(x-4)^2}\]\[=\frac{-x^2 + 5x + 2}{(x-4)^2}\]\[=\frac{-(x^2 - 5x - 2)}{(x-4)^2}\]\[=-\frac{x^2 - 5x - 2}{(x-4)^2}\]

OpenStudy (anonymous):

thanks allot for the help!